OpenStudy (anonymous):

Which shows the expressions in the order they would appear on a number line from least to greatest?

OpenStudy (anonymous):

4^3 the S.R. of 40 the S.R. of 13 the S.R. of 7 14/9

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@shrutipande9

OpenStudy (anonymous):

4^3, the S.R. of 40, the S.R. of 13, the S.R. of 7, 14/9.

OpenStudy (anonymous):

@AngelWilliams16

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Which shows the expressions in the order they would appear on a number line from least to greatest? 4^3, the S.R. of 40, the S.R. of 13, the S.R. of 7, 14/9.

OpenStudy (igreen):

Simplify them to whole numbers/decimals.

OpenStudy (igreen):

\(4^3 = 64\) \(\sqrt{40} \approx 6.32\) \(\sqrt{13} \approx 3.61\) \(\sqrt7 \approx 2.65\) \(\dfrac{14}{9} \approx 1.56\)

OpenStudy (anonymous):

ok.

OpenStudy (igreen):

So in order from least to greatest it's going to be: \(\dfrac{14}{9}, \sqrt7, \sqrt{13}, \sqrt{40}, 4^3\)

OpenStudy (anonymous):

ty :P

OpenStudy (igreen):

Np

OpenStudy (anonymous):

tht is an answer

OpenStudy (anonymous):

Which rational number equals 0.1 with bar over 1

OpenStudy (anonymous):

i can remember the name xd

OpenStudy (anonymous):

bruh.

OpenStudy (anonymous):

for this,

OpenStudy (anonymous):

The amounts of e-waste generated in a region during two years were 14,200,000 tons and 10,700,000 tons. What was the total e-waste generated over these two years in the region, expressed in scientific notation?

OpenStudy (anonymous):

i know i have to add the convert into scientific notation

OpenStudy (anonymous):

but idk where to put the decimal point

OpenStudy (anonymous):

i got it nvm

OpenStudy (igreen):

Lol

OpenStudy (anonymous):

2.49 x 10^7

OpenStudy (anonymous):

one last question

OpenStudy (anonymous):

what expression is equivalent to,

OpenStudy (igreen):

Just move the decimal place 7 times to the right: 24900000

OpenStudy (anonymous):

3^2 x 3^-5 what would be equivalent to this?

OpenStudy (igreen):

When you multiply exponents with the like base, you add the exponents and you keep the base. \(Base\rightarrow 5^{2\leftarrow Exponent}\) Add the exponents: 2 + -5 = -3 Keep the base, 3. So we have \(3^{-3}\).

OpenStudy (anonymous):

nice :P

OpenStudy (anonymous):

and ty

OpenStudy (anonymous):

pass :D

OpenStudy (igreen):

:P